OpenStudy (anonymous):

ugh stuck for some reason 1/10 + 2/5 = ??? same donomanator? anyone

OpenStudy (anonymous):

look! 10*5

OpenStudy (anonymous):

then 50/10 = 5

OpenStudy (anonymous):

then 5*1 = 5

OpenStudy (anonymous):

Now: 50/5 = 10

OpenStudy (anonymous):

and 10 * 2= 20

OpenStudy (anonymous):

so : 5+20 = 25

OpenStudy (anonymous):

25/50

OpenStudy (anonymous):

\[\frac{1}{10}+\frac{2}{5}*\frac{2}{2}=\frac{1}{10}+\frac{4}{10}=\frac{5}{10}=\frac{1}{2}\]

OpenStudy (anonymous):